OpenStudy (anonymous):

Can someone help me setup this rate of work problem? It takes 7 hours for Isabella to rake leaves by herself, but her brother Matthew can work three times as fast. If they work together how long will it take them to rake leaves? I'm given a chart with boxes for Isabella and Matthew. They are labeled: Rate Time Work Done

OpenStudy (anonymous):

So if matthew can work three times as fast how long will it take him to rake leaves alone. also is this the whole problem because there should be some information on the amount of leaves

OpenStudy (anonymous):

This is the whole problem.

OpenStudy (anonymous):

Isabella takes 7 hours to rake leaves. Her brother takes 7/3 hours to rake same no. of leaves. when they work together their work of one hour\[=\frac{ 1 }{ 7 }+\frac{ 3 }{ 7 }=\frac{ 1+3 }{ 7 }=\frac{ 4 }{ 7 }\] Together they can rake leaves in 7/4 hours.

OpenStudy (anonymous):

Thanks @surjithayer

OpenStudy (anonymous):

How would i set it up in this chart?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

here rate=1/7 time=1 hour work done =1/7 th